the sum of two numbers is 84.if the difference between half of the first number and one third of the second number is 2,find the two numbers
Answer by VFBundy(438)

First number = x
Second number = y

x + y = 84
x/2 - y/3 = 2

Multiply the second equation by 3:

3(x/2 - y/3) = 3(2)

3x/2 - y = 6

Now, here are your two equations:

x + y = 84
3x/2 - y = 6

Add the two equations together:

5x/2 = 90

Solve for x:

5x/2 = 90

5x = 180

x = 36

Now that we know x = 36, we can plug this into any of the previous equations to find y. Use x + y = 84:

x + y = 84

36 + y = 84

y = 48

First number = x = 36
Second number = y = 48